# SymTFT out of equilibrium: from time crystals to braided drives and Floquet codes

### 28 Dec 2023 v1

@@ Abstract
Symmetry Topological Field Theory (SymTFT) is a framework to capture universal features of quantum many-body systems by viewing them as a boundary of topological order in one higher dimension. This yielded numerous insights in static low-energy settings. Here we study what SymTFT can tell about nonequilibrium, focusing on one-dimensional (1D) driven systems and their 2D SymTFT. In driven settings, boundary conditions (BCs) can be dynamical and can apply both spatially and temporally. We show how this enters SymTFT via topological operators, which we then use to uncover several new results for 1D dynamics. These include revealing time crystals (TCs) as systems with symmetry-twisted temporal BCs, finding robust bulk TC features in phases thought to be only boundary TCs, Floquet dualities, or identifying Floquet codes as space-time duals to systems with duality-twisted spatial BCs. We also show how, by making duality-twisted BCs dynamical, non-Abelian braiding of duality defects can enter SymTFT, leading to effects such as the exact pumping of symmetry charges between a system and its BCs. We illustrate our ideas for $\mathbb{Z}_{2}$-symmetric 1D systems, but our construction applies for any finite Abelian symmetry.
@@
Recent years saw the development of deep generalizations of symmetry Nussinov and Ortiz (2009a, b); Cobanera et al. (2010, 2011); Kapustin and Seiberg (2014); Gaiotto et al. (2015); Yoshida (2016); Lake ; McGreevy (2023); Schafer-Nameki ; Bhardwaj et al. (2024). Among these is the insight that upon viewing a system as a boundary of topological order (TO) Kitaev (2003, 2006); Levin and Wen (2005); Nayak et al. (2008) in one higher dimension, one can study its symmetries, their extensions to categorical symmetries involving dualities and “dual” symmetries, and various twisted boundary conditions (BCs) in a “theory independent” manner, i.e., insensitive to many microscopic details Severa (2002); Fröhlich et al. (2004, 2007); Kitaev and Kong (2012); Nussinov et al. (2012); Cobanera et al. (2013); Kong et al. (2017); Thorngren and Wang ; Albert et al. ; Freed and Teleman ; Gaiotto and Kulp (2021); Ho et al. (2015); Aasen et al. (2016, ); Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Chatterjee and Wen (2023); Moradi et al. (2023); Ji and Wen (2020); Lichtman et al. (2021); Freed et al. ; Lootens et al. (2021, 2022). The TO here is sometimes called “symmetry topological field theory” (SymTFT) Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Chatterjee and Wen (2023); Moradi et al. (2023).
@@
This uncovered unitary braided tensor categories as structures unifying topological defects describing dualities and symmetries Freed and Teleman ; Gaiotto and Kulp (2021); Ho et al. (2015); Aasen et al. (2016, ); Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Chatterjee and Wen (2023); Ji and Wen (2020); Lichtman et al. (2021); Moradi et al. (2023); Freed et al. ; Lootens et al. (2021, 2022). The non-Abelian fusion in these categories underlie viewing them as “non-invertible” symmetries. The generality of SymTFT notwithstanding, it has thus far been used only in static low-energy settings, yielding insights including classifications of gapped phases, dualities, or critical properties Freed and Teleman ; Gaiotto and Kulp (2021); Ho et al. (2015); Aasen et al. (2016, ); Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Chatterjee and Wen (2023); Ji and Wen (2020); Lichtman et al. (2021); Moradi et al. (2023); Freed et al. ; Lootens et al. (2021, 2022, 2023); Bhardwaj et al. ; Bhardwaj et al. ; Pace ; Wen and Potter ; Huang and Cheng .
@@
What can SymTFT tell about nonequilibrium? Here we study this, focusing on one-dimensional (1D) driven systems and their 2D SymTFT. We take the former many-body localized (MBL) Fleishman and Anderson (1980); Gornyi et al. (2005); Basko et al. (2006); Huse et al. (2013); Serbyn et al. (2013); Huse et al. (2014); Chandran et al. (2015); Ros et al. (2015); Wahl et al. (2017); Goihl et al. (2018); Nandkishore and Huse (2015); Altman and Vosk (2015); Abanin and Papić (2017); Alet and Laflorencie (2018); Imbrie et al. (2017); this allows for genuinely dynamical phases, e.g., time crystals (TCs) Wilczek (2012); Shapere and Wilczek (2012); Watanabe and Oshikawa (2015); Khemani et al. (2016); Else et al. (2016); von Keyserlingk and Sondhi (2016a, b); von Keyserlingk et al. (2016); Khemani et al. (2017, ); Sacha and Zakrzewski (2018); Else et al. (2020).
@@
Driven settings can involve both spatial and temporal BCs, and BCs can be dynamical. This adds novel features to SymTFT, while yielding new insights and predictions in 1D, including revealing TCs as systems with symmetry-twisted temporal BCs, robust bulk TC features in phases thought to be only boundary TCs, Floquet dualities, or Floquet codes (FCs, of great interest in quantum computing Hastings and Haah (2021); Vuillot ; Davydova et al. (2023); Kesselring et al. ; Vu et al. ; Sullivan et al. (2023); Aasen et al. ; Ellison et al. ; Davydova et al. ; Dua et al. ) as space-time duals to systems with duality-twisted spatial BCs. One of our most striking results is that, by making duality-twisted BCs dynamical, non-Abelian braiding—not only the aforementioned fusion—of duality defects Bombin (2010); Barkeshli and Qi (2012); You and Wen (2012); Barkeshli et al. (2013a, 2019) can enter SymTFT, with remarkable consequences such as the exact pumping of symmetry charges between a system and its BCs.
@@
Hilbert space, symmetries, order parameters: We first summarize the (condensed matter) SymTFT construction, bridging between Refs. Ji and Wen, 2020; Lichtman et al., 2021; Chatterjee and Wen, 2023; Moradi et al., 2023 by our approach to order parameters, degeneracies, and BCs SM . We focus on 1D spin chains with closed BCs and $L$-site Hilbert space $\mathcal{H}_{\text{dyn}}=\mathcal{H}_{1}^{\otimes L}$; for $G=\mathbb{Z}_{2}$, $\mathcal{H}_{1}=\text{span}(\{\ket{0},\ket{1}\})$ SM . The global symmetry acts as $V_{g}=\prod_{k=1}^{L}v_{g}^{(k)}$, $g\in G$, with onsite unitary $v_{g}^{(k)}$ at site $k$. The “patch” symmetry Ji and Wen (2020), for interval (“patch”) $[i,j]$ is $V_{g}^{(ij)}=\prod_{k=i}^{j}v_{g}^{(k)}$.
@@
For the 2D setup to reduce to 1D, we require that away from $B_{\text{dyn}}$ there be a gap much larger than $B_{\text{dyn}}$ energy scales. Thus, we identify $\mathcal{H}_{\text{dyn}}$ as the subspace with zero anyons away from $B_{\text{dyn}}$. The 1D algebra is then the “boundary algebra” (BA): those of the TO’s unitary string operators $W^{(a)}$ (more precisely, their equivalence classes under endpoint-preserving deformations) that may create anyons $a$ at and only at $B_{\text{dyn}}$. Anyons are the endpoints of $W^{(a)}$. Thus, the BA includes open strings $W_{ij}^{(a)}$ ending on an $a$-$\bar{a}$ pair at $i,j\in B_{\text{dyn}}$, cf. Fig. 1 Wde . By the TO’s fusion and braiding properties, respectively Ji and Wen (2020); Lichtman et al. (2021); Chatterjee and Wen (2023); Moradi et al. (2023),
@@
To relate Eq. (1) to the 1D algebra, we first identify $W_{i-1,j}^{(g,1)}=V_{g}^{(ij)}$. For $G=\mathbb{Z}_{2}$ this is $W_{i-1,j}^{(m)}=\prod_{i\leq k\leq j}X_{k}$, with $X_{k}$, $Y_{k}$, $Z_{k}$ the Pauli operators on site $k$. The global symmetry is $\smash{\overline{W}}^{(g,1)}=V_{g}$, where $\smash{\overline{W}}^{(g,1)}$ encircles $B_{\text{dyn}}$, cf. Fig. 1. For $G=\mathbb{Z}_{2}$, $\smash{\overline{W}}^{(m)}=\prod_{k=1}^{L}X_{k}$.
@@
Next, we implement local (on-site) order parameters $O_{k}^{(\alpha)}$, obeying $V_{g}^{(ij)}O_{k}^{(\alpha)}V_{g}^{(ij)\dagger}=\chi_{\alpha}(g)O_{k}^{(\alpha)}$ for $i\leq k\leq j$. By $W_{i-1,j}^{(g,1)}=V_{g}^{(ij)}$, we get the same relation if $O_{k}^{(\alpha)}=W_{k}^{(1,\alpha)}$, a string with a single endpoint $k\in B_{\text{dyn}}$. To ensure that the other endpoint creates no anyons, we place the 2D TO on a cylinder and emanate $W_{k}^{(1,\alpha)}$ from the other, gapped, “reference” boundary $B_{\text{ref}}$, with Lagrangian subgroup $\mathcal{A}_{\text{ref}}=\langle\{(1,\alpha)|\alpha\in\text{Rep}\,G\}\rangle$, cf. Fig. 1. For $G=\mathbb{Z}_{2}$, we have $\mathcal{A}_{\text{ref}}=\langle e\rangle$, $W_{k}^{(e)}=Z_{k}$.
@@
While $W_{ij}^{(g,1)}$ and $W_{k}^{(1,\alpha)}$ express standard 1D concepts, we must comment on their “electric-magnetic dual” $W_{ij}^{(1,\alpha)}$ and $W_{\tilde{k}}^{(g,1)}$ Wde . The $W_{ij}^{(1,\alpha)}$ describe the “dual symmetry” $\tilde{G}$, the one broken in a disordered phase Ji and Wen (2020). By incorporating both $G$ and $\tilde{G}$, Eq. (1) expresses the categorical symmetries in 1D Ji and Wen (2020). For $G=\mathbb{Z}_{2}$, we denote $\tilde{G}=\tilde{\mathbb{Z}}_{2}$, and by $W_{i}^{(1,\alpha)\dagger}W_{j}^{(1,\alpha)}=W_{ij}^{(1,\alpha)}$ (via deformations), we have $W_{ij}^{(e)}=Z_{i}Z_{j}$. For the “dual order parameter” $W_{\tilde{k}}^{(g,1)}$, we must discuss twisted BCs. This is what we do next, via “duality defects” on the cylinder.
@@
Twisted spatial boundary BCs: TO comes with anyonic symmetries $\sigma$ Kitaev (2006); Bombin (2010); Barkeshli and Qi (2012); You and Wen (2012); Barkeshli et al. (2013a, 2019): automorphisms $a^{\prime}=\sigma(a)$ preserving fusion and braiding properties. They may furnish domain walls $D^{(\sigma)}$ such that string $W^{(a)}$ must continue as $W^{[\sigma(a)]}$ across $D^{(\sigma)}$ to avoid creating an anyon at $D^{(\sigma)}$, cf. Fig. 1. A finite-length $D^{(\sigma)}$ ends on “twist defects” Bombin (2010); Barkeshli and Qi (2012); You and Wen (2012); Barkeshli et al. (2013a, 2019). Automorphisms include electric-magnetic dualities, $\sigma(g)\in\text{Rep}\,G$, $\forall g\in G$, $\sigma^{2}=1$. We mostly focus on these henceforth. For $G=\mathbb{Z}_{2}$ this is $\sigma(e)=m$.
@@
With $D^{(\sigma)}$ we now $g$-twist BCs. (We return to duality-twisting later.) To this end, we add a short wall $D_{\text{ref}}^{(\sigma)}$ near $B_{\text{ref}}$, cf. Fig. 1; crossing this, $(1,\alpha)$ emanating at $B_{\text{\text{ref}}}$ yields $W_{\tilde{j}}^{(g,1)}$ with $g=\sigma(\alpha)$, i.e., an unpaired $(g,1)$ anyon at $B_{\text{dyn}}$. This $g$-twists BCs via the $G$-Aharonov-Bohm effect: $\smash{\overline{W}}^{(1,\alpha)}$, bringing a charge $(1,\alpha)$ around the system, detects the flux $g$ by $\smash{\overline{W}}^{(1,\alpha)}\!\ket{\psi^{\prime}}=\chi_{\alpha}(g)\ket{\psi^{\prime}}$, where |$\ket{\psi^{\prime}}=W_{\tilde{j}}^{(g,1)}\!\ket{\psi}$ and $\smash{\overline{W}}^{(1,\alpha)}\ket{\psi}=\ket{\psi}$ SM .
@@
Figure 1: Embedding the 1D system as boundary $B_{\text{dyn}}$ of a 2D topological order on a cylinder (dashed sides are identified). Boundary $B_{\text{ref}}$ condenses $(1,\alpha)$ anyons. $W^{(1,\alpha)}$ lines are shown in red, $W^{(g,1)}$ in blue. The $W_{i}^{(1,\alpha)}$ serve as local order parameters in 1D; $\smash{\overline{W}}^{(g,1)}$ are global symmetries. Also shown are $W^{(a)}_{j-2,j-1}$ for $a\in\{(1,\alpha),(g,1)\}$, as is $W^{(1,\alpha)}_{l,l+1}$, crossing domain wall $D_{\text{dyn}}^{(\sigma)}$. $D_{\text{ref}}^{(\sigma)}$ enables $g$-twisting boundary conditions via $W_{\tilde{j}}^{(g,1)}$. Crosses represent twist defects.
@@
For $G=\mathbb{Z}_{2}$, $W_{\tilde{j}}^{(m)}W_{k,k+1}^{(e)}W_{\tilde{j}}^{(m)\dagger}=(-1)^{\delta_{jk}}W_{k,k+1}^{(e)}$; thus $W_{\tilde{j}}^{(m)}$ flips $Z_{k}Z_{k+1}$ only for $k=j$. No $X_{k}$ product does this. Thus, we gauge via $Z_{j}Z_{j+1}\to Z_{j}\tilde{Z}_{\tilde{j}}Z_{j+1}\equiv W_{j,j+1}^{(e)}$, yielding $W_{\tilde{j}}^{(m)}=\tilde{X}_{\tilde{j}}$, with link Paulis $\tilde{Z}_{\tilde{j}}$, $\tilde{X}_{\tilde{j}}$. In a gauge with just one dynamic link, i.e., $W_{k,k+1}^{(e)}=Z_{k}Z_{k+1}$ ($k\neq j$), we have $\smash{\overline{W}}^{(e)}=\prod_{j}W_{j,j+1}^{(e)}=\tilde{Z}_{\tilde{j}}$.
@@
$\pi$SG SymTFT: We first illustrate this for the paradigmatic $\mathbb{Z}_{2}$ TC, the $\pi$SG Khemani et al. (2016); Else et al. (2016); von Keyserlingk and Sondhi (2016a, b); von Keyserlingk et al. (2016); Khemani et al. (2017, ); Sacha and Zakrzewski (2018); Else et al. (2020). This arises for $b=m$ and $\mathcal{A}=\langle e\rangle$. We start with the zero-correlation-length (or fixed-point, FP) limit, $H_{1}=0$. Here $W_{k,k+1}^{(e)}$ are local integrals of motion (LIOMs) Serbyn et al. (2013); Huse et al. (2014); Chandran et al. (2015); Ros et al. (2015); Wahl et al. (2017); Goihl et al. (2018); Nandkishore and Huse (2015); Altman and Vosk (2015); Abanin and Papić (2017); Alet and Laflorencie (2018); Imbrie et al. (2017) and the eigenstates of $U_{\text{F}}^{(m,\langle e\rangle)}$ are |$\ket{n}=\ket{\{d_{k}\},p}$, with $d_{k},p\in\{-1,1\}$ where $W_{k,k+1}^{(e)}\ket{n}=d_{k}\ket{n}$, $\smash{\overline{W}}^{(m)}\ket{n}=p\ket{n}$. These are SG eigenstates, displaying long-ranged order-parameter correlations $C_{n;ij}=\bra{n}W_{ij}^{(e)}\ket{n}\neq 0$. (We henceforth write $C_{n;ij}=C_{ij}$ for brevity.)
Furthermore, by $\smash{\overline{W}}^{(m)\dagger}W_{i}^{(e)}\smash{\overline{W}}^{(m)}=e^{-2i\theta_{em}}W_{i}^{(e)}$ and $H_{0}^{(\langle e\rangle)}W_{i}^{(e)}=W_{i}^{(e)}H_{0}^{(\langle e\rangle)}$ (for $H_{1}=0$), we have $W_{i}^{(e)}(t)=[U_{\text{F}}^{(m,\langle e\rangle)\dagger}]^{t}W_{i}^{(e)}[U_{\text{F}}^{(m,\langle e\rangle)}]^{t}=e^{-2it\theta_{em}}W_{i}^{(e)}$ ($t\in\mathbb{Z}$) hence $C_{ij}(t)=\bra{n}W_{i}^{(e)}(t)W_{j}^{(e)}\ket{n}=(-1)^{t}C_{ij}$. This period doubled long-ranged order is the spatio-temporal order defining TCs von Keyserlingk et al. (2016); Khemani et al. (2017, ). It also comes with spectral multiplets: if $U_{\text{F}}^{(m,\langle e\rangle)}\ket{n}=e^{-i\varepsilon_{n}}\ket{n}$, then, for $H_{1}=0$, $U_{\text{F}}^{(m,\langle e\rangle)}W_{i}^{(e)}\ket{n}=e^{2i\theta_{em}}e^{-i\varepsilon_{n}}W_{i}^{(e)}\ket{n}$; by $2\theta_{em}=\pi$, this is the familiar $\pi$ spectral pairing von Keyserlingk et al. (2016); Khemani et al. (2017, ).

### 28 Feb 2024 v2

@@ Abstract
Symmetry Topological Field Theory (SymTFT) is a framework to capture universal features of quantum many-body systems by viewing them as a boundary of topological order in one higher dimension. This has yielded numerous insights in static low-energy settings. Here we study what SymTFT can reveal about nonequilibrium, focusing on one-dimensional (1D) periodically driven systems and their 2D SymTFTs. In driven settings, boundary conditions (BCs) can be dynamical and can apply both spatially and temporally. We show how this enters SymTFT via topological operators, which we then use to uncover several new results. These include revealing time crystals (TCs) as systems with symmetry-twisted temporal BCs, robust bulk “dual TCs” in phases thought to be only boundary TCs, generating drive dualities, or identifying 2D Floquet codes as space-time duals to 1D systems with duality-twisted spatial BCs. We also show how, by making duality-twisted BCs dynamical, non-Abelian braiding of duality defects can enter SymTFT, leading to effects such as the exact pumping of symmetry charges between a system and its BCs. We illustrate our ideas for $\mathbb{Z}_{2}$-symmetric 1D systems, but our construction applies for any finite Abelian symmetry.
@@
Recent years saw the development of deep generalizations of symmetry Nussinov and Ortiz (2009a, b); Cobanera et al. (2010, 2011); Kapustin and Seiberg (2014); Gaiotto et al. (2015); Yoshida (2016); Lake ; McGreevy (2023); Schafer-Nameki ; Bhardwaj et al. (2024). Among these is the insight that upon viewing a system as a boundary of topological order (TO) Kitaev (2003, 2006); Levin and Wen (2005); Nayak et al. (2008) in one higher dimension, one can study its symmetries, their extensions to categorical symmetries involving dualities and “dual” symmetries, and various twisted boundary conditions (BCs) in a “theory independent” manner, i.e., insensitive to many microscopic details Severa (2002); Fröhlich et al. (2004, 2007); Kitaev and Kong (2012); Nussinov et al. (2012); Cobanera et al. (2013); Kong et al. (2017); Thorngren and Wang ; Albert et al. ; Freed and Teleman ; Gaiotto and Kulp (2021); Ho et al. (2015); Aasen et al. (2016, ); Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Ji and Wen (2020); Lichtman et al. (2021); Chatterjee and Wen (2023a); Moradi et al. (2023); Freed et al. ; Lootens et al. (2021, 2022). The TO here is sometimes called “symmetry topological field theory” (SymTFT) Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Chatterjee and Wen (2023a); Moradi et al. (2023).
@@
This uncovered unitary braided tensor categories as structures unifying topological defects describing dualities and symmetries Freed and Teleman ; Gaiotto and Kulp (2021); Ho et al. (2015); Aasen et al. (2016, ); Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Chatterjee and Wen (2023a); Ji and Wen (2020); Lichtman et al. (2021); Moradi et al. (2023); Freed et al. ; Lootens et al. (2021, 2022). The non-Abelian fusion in these categories render them into “non-invertible” symmetries.
The generality of SymTFT notwithstanding, it has thus far been used only in static low-energy settings, yielding insights including classifications of gapped phases, dualities, or critical properties Freed and Teleman ; Gaiotto and Kulp (2021); Ho et al. (2015); Aasen et al. (2016, ); Apruzzi et al. ; Kaidi et al. ; Bhardwaj and Schafer-Nameki ; Chatterjee and Wen (2023a); Ji and Wen (2020); Lichtman et al. (2021); Moradi et al. (2023); Freed et al. ; Lootens et al. (2021, 2022, 2023); Bhardwaj et al. ; Bhardwaj et al. ; Pace ; Chatterjee and Wen (2023b); Wen and Potter ; Huang and Cheng ; Bhardwaj et al. .
@@
What can SymTFT reveal about nonequilibrium? Here we study this, focusing on one-dimensional (1D) Floquet systems and their 2D SymTFTs. We take the former many-body localized (MBL) Fleishman and Anderson (1980); Gornyi et al. (2005); Basko et al. (2006); Huse et al. (2013); Serbyn et al. (2013); Huse et al. (2014); Chandran et al. (2015); Ros et al. (2015); Wahl et al. (2017); Goihl et al. (2018); Nandkishore and Huse (2015); Altman and Vosk (2015); Abanin and Papić (2017); Alet and Laflorencie (2018); Imbrie et al. (2017); this allows for genuinely dynamical phases, e.g., time crystals (TCs) Wilczek (2012); Shapere and Wilczek (2012); Watanabe and Oshikawa (2015); Khemani et al. (2016); Else et al. (2016); von Keyserlingk and Sondhi (2016a, b); von Keyserlingk et al. (2016); Khemani et al. (2017, ); Sacha and Zakrzewski (2018); Else et al. (2020).
@@
Driven settings can involve both spatial and temporal BCs, and BCs can be dynamical. As we shall show, this adds novel features to SymTFT which, in turn, yield new insights in 1D that reveal, e.g., TCs as featuring symmetry-twisted temporal BCs, bulk “dual TCs” in phases thought to be only boundary TCs, or dynamical dualities. We also show how space-time duality of BCs, when applied to duality-twisted BCs, yields 2D Floquet codes (FCs) Hastings and Haah (2021); Vuillot ; Davydova et al. (2023); Kesselring et al. ; Vu et al. ; Sullivan et al. (2023); Aasen et al. ; Ellison et al. ; Davydova et al. ; Dua et al. , thus unifying TCs and FCs via SymTFT. Furthermore we show that, by making duality-twisted BCs dynamical, non-Abelian braiding—not only fusion—of duality defects Bombin (2010); Barkeshli and Qi (2012); You and Wen (2012); Barkeshli et al. (2013a, 2019) can enter SymTFT, with remarkable consequences such as the exact pumping of symmetry charges between a system and its BCs.
@@
Hilbert space, symmetries, order parameters: We first summarize the (condensed matter) SymTFT construction, bridging between Refs. Ji and Wen, 2020; Lichtman et al., 2021; Chatterjee and Wen, 2023a; Moradi et al., 2023 by our approach to order parameters, degeneracies, and BCs SM . We focus on 1D spin chains with closed BCs and $L$-site Hilbert space $\mathcal{H}_{\text{dyn}}=\mathcal{H}_{1}^{\otimes L}$; for $G=\mathbb{Z}_{2}$, $\mathcal{H}_{1}=\text{span}(\{\ket{0},\ket{1}\})$ SM . The global symmetry acts as $V_{g}=\prod_{k=1}^{L}v_{g}^{(k)}$, $g\in G$, with onsite unitary $v_{g}^{(k)}$ at site $k$. The “patch” symmetry Ji and Wen (2020), for interval (“patch”) $[i,j]$ is $V_{g}^{(ij)}=\prod_{k=i}^{j}v_{g}^{(k)}$.
@@
For the 2D setup to reduce to 1D, we require that away from $B_{\text{dyn}}$ there be a gap much larger than $B_{\text{dyn}}$ energy scales. Thus, we identify $\mathcal{H}_{\text{dyn}}$ as the subspace with zero anyons away from $B_{\text{dyn}}$. The 1D algebra is then the “boundary algebra” (BA): that of the TO’s unitary string operators $W^{(a)}$ (more precisely, their equivalence classes under endpoint-preserving deformations) that may create anyons $a$ at and only at $B_{\text{dyn}}$. Anyons are the endpoints of $W^{(a)}$. Thus, the BA includes open strings $W_{ij}^{(a)}$ ending on an $a$-$\bar{a}$ pair at $i,j\in B_{\text{dyn}}$, cf. Fig. 1 Wde . By the TO’s fusion and braiding properties, respectively Ji and Wen (2020); Lichtman et al. (2021); Chatterjee and Wen (2023a); Moradi et al. (2023),
@@
To relate Eq. (1) to the 1D algebra, we first identify $W_{i-1,j}^{(g,1)}\equiv V_{g}^{(ij)}$ Wde . For $G=\mathbb{Z}_{2}$ this is $W_{i-1,j}^{(m)}=\prod_{i\leq k\leq j}X_{k}$, with $X_{k}$, $Y_{k}$, $Z_{k}$ the Pauli operators on site $k$. The global symmetry is $\smash{\overline{W}}^{(g,1)}=V_{g}$, where $\smash{\overline{W}}^{(g,1)}$ encircles $B_{\text{dyn}}$, cf. Fig. 1. For $G=\mathbb{Z}_{2}$, $\smash{\overline{W}}^{(m)}=\prod_{k=1}^{L}X_{k}$.
@@
Next, we implement local (on-site) order parameters $O_{k}^{(\alpha)}$, obeying $V_{g}^{(ij)}O_{k}^{(\alpha)}V_{g}^{(ij)\dagger}=\chi_{\alpha}(g)O_{k}^{(\alpha)}$ for $i\leq k\leq j$. By $W_{i-1,j}^{(g,1)}=V_{g}^{(ij)}$, we get the same relation if $O_{k}^{(\alpha)}=W_{k}^{(1,\alpha)}$, a string with a single endpoint $k\in B_{\text{dyn}}$. To ensure that the other endpoint creates no anyons, we place the 2D TO on a cylinder and emanate $W_{k}^{(1,\alpha)}$ from the other, gapped, “reference” boundary $B_{\text{ref}}$, with Lagrangian subgroup $\mathcal{A}_{\text{ref}}=\langle\{(1,\alpha)|\alpha\in\text{Rep}\,G\}\rangle$, cf. Fig. 1. For $G=\mathbb{Z}_{2}$, we have $\mathcal{A}_{\text{ref}}=\langle e\rangle$, $W_{k}^{(e)}\equiv Z_{k}$.
@@
While $W_{ij}^{(g,1)}$ and $W_{k}^{(1,\alpha)}$ express standard 1D concepts, we must comment on their “electric-magnetic dual” $W_{ij}^{(1,\alpha)}$ and $W_{k}^{(g,1)}$ Wde . The $W_{ij}^{(1,\alpha)}$ describe the “dual symmetry” $\tilde{G}$, the one broken in a disordered phase Ji and Wen (2020). By incorporating both $G$ and $\tilde{G}$, Eq. (1) expresses the categorical symmetries in 1D Ji and Wen (2020). For $G=\mathbb{Z}_{2}$, we denote $\tilde{G}=\tilde{\mathbb{Z}}_{2}$, and by $W_{i}^{(1,\alpha)\dagger}W_{j}^{(1,\alpha)}=W_{ij}^{(1,\alpha)}$ (via deformations), we have $W_{ij}^{(e)}=Z_{i}Z_{j}$. For the “dual order parameter” $W_{k}^{(g,1)}$, we must discuss twisted BCs. This is what we do next, via “duality defects” on the cylinder.
@@
Twisted spatial BCs: TO comes with anyonic symmetries $\sigma$ Kitaev (2006); Bombin (2010); Barkeshli and Qi (2012); You and Wen (2012); Barkeshli et al. (2013a, 2019): automorphisms $a^{\prime}=\sigma(a)$ preserving fusion and braiding properties. They may furnish domain walls $D^{(\sigma)}$ such that string $W^{(a)}$ must continue as $W^{[\sigma(a)]}$ across $D^{(\sigma)}$ to avoid creating an anyon at $D^{(\sigma)}$, cf. Fig. 1. A finite-length $D^{(\sigma)}$ ends on “twist defects” Bombin (2010); Barkeshli and Qi (2012); You and Wen (2012); Barkeshli et al. (2013a, 2019). Automorphisms include electric-magnetic dualities, $\sigma(g)\in\text{Rep}\,G$, $\forall g\in G$, $\sigma^{2}=1$. We mostly focus on these henceforth. For $G=\mathbb{Z}_{2}$ this is $\sigma(e)=m$.
@@
With $D^{(\sigma)}$ we now $g$-twist BCs. (We return to duality-twisting later.) To this end, we add a short wall $D_{\text{ref}}^{(\sigma)}$ near $B_{\text{ref}}$, cf. Fig. 1; crossing this, $(1,\alpha)$ emanating at $B_{\text{\text{ref}}}$ yields $W_{j}^{(g,1)}$ with $g=\sigma(\alpha)$, i.e., an unpaired $(g,1)$ anyon at $B_{\text{dyn}}$ Wde . This $g$-twists BCs via the $G$-Aharonov-Bohm effect: $\smash{\overline{W}}^{(1,\alpha)}$, bringing a charge $(1,\alpha)$ around the system, detects the flux $g$ by $\smash{\overline{W}}^{(1,\alpha)}\!\ket{\psi^{\prime}}=\chi_{\alpha}(g)\ket{\psi^{\prime}}$, where |$\ket{\psi^{\prime}}=W_{j}^{(g,1)}\!\ket{\psi}$ and $\smash{\overline{W}}^{(1,\alpha)}\ket{\psi}=\ket{\psi}$ SM .
@@
Figure 1: Embedding the 1D system as boundary $B_{\text{dyn}}$ of a 2D TO on a cylinder (dashed sides are identified). Boundary $B_{\text{ref}}$ condenses $(1,\alpha)$ anyons. $W^{(1,\alpha)}$ lines are shown in red, $W^{(g,1)}$ in blue. The $W_{i}^{(1,\alpha)}$ serve as local order parameters in 1D; $\smash{\overline{W}}^{(g,1)}$ are global symmetries. Also shown are $W^{(a)}_{j-2,j-1}$ for $a\in\{(1,\alpha),(g,1)\}$, as is $W^{(1,\alpha)}_{l,l+1}$, crossing domain wall $D_{\text{dyn}}^{(\sigma)}$. $D_{\text{ref}}^{(\sigma)}$ enables $g$-twisting boundary conditions via $W_{j}^{(g,1)}$ (ending on link $\langle j,j+1\rangle$Wde . Crosses represent twist defects.
@@
For $G=\mathbb{Z}_{2}$, $W_{j}^{(m)}W_{k,k+1}^{(e)}W_{j}^{(m)\dagger}=(-1)^{\delta_{jk}}W_{k,k+1}^{(e)}$; thus $W_{j}^{(m)}$ flips $Z_{k}Z_{k+1}$ only for $k=j$. No $X_{k}$ product does this. Thus, we gauge via $Z_{j}Z_{j+1}\to Z_{j}\tilde{Z}_{j}Z_{j+1}\equiv W_{j,j+1}^{(e)}$, yielding $W_{j}^{(m)}\equiv\tilde{X}_{j}$, with link Paulis $\tilde{Z}_{j}$, $\tilde{X}_{j}$. In a gauge with just one dynamic link, i.e., $W_{k,k+1}^{(e)}=Z_{k}Z_{k+1}$ ($k\neq j$), we have $\smash{\overline{W}}^{(e)}=\prod_{j=1}^{L}W_{j,j+1}^{(e)}=\tilde{Z}_{j}$ detecting untwisted (periodic) and $m$-twisted (antiperiodic) BCs SM .
@@
TC SymTFT: We first illustrate this for the paradigmatic $\mathbb{Z}_{2}$ TC, the $\pi$SG Khemani et al. (2016); Else et al. (2016); von Keyserlingk and Sondhi (2016a, b); von Keyserlingk et al. (2016); Khemani et al. (2017, ); Sacha and Zakrzewski (2018); Else et al. (2020). This arises for $b=m$ and $\mathcal{A}=\langle e\rangle$. We start with the zero-correlation-length (or fixed-point, FP) limit, $H_{1}=0$. Here $W_{k,k+1}^{(e)}$ are local integrals of motion (LIOMs) Serbyn et al. (2013); Huse et al. (2014); Chandran et al. (2015); Ros et al. (2015); Wahl et al. (2017); Goihl et al. (2018); Nandkishore and Huse (2015); Altman and Vosk (2015); Abanin and Papić (2017); Alet and Laflorencie (2018); Imbrie et al. (2017) and the eigenstates of $U_{\text{F}}^{(m,\langle e\rangle)}$ are |$\ket{n}=\ket{\{d_{k}\},p}$, with $d_{k},p\in\{-1,1\}$ where $W_{k,k+1}^{(e)}\ket{n}=d_{k}\ket{n}$, $\smash{\overline{W}}^{(m)}\ket{n}=p\ket{n}$. These are SG eigenstates, displaying long-ranged order-parameter correlations $C_{n;ij}=\bra{n}W_{ij}^{(e)}\ket{n}\neq 0$. (We henceforth write $C_{n;ij}=C_{ij}$ for brevity.)
Furthermore, by $\smash{\overline{W}}^{(m)\dagger}W_{i}^{(e)}\smash{\overline{W}}^{(m)}=e^{-2i\theta_{em}}W_{i}^{(e)}$ and $H_{0}^{(\langle e\rangle)}W_{i}^{(e)}=W_{i}^{(e)}H_{0}^{(\langle e\rangle)}$, at the FP we have $W_{i}^{(e)}(t)=[U_{\text{F}}^{(m,\langle e\rangle)\dagger}]^{t}W_{i}^{(e)}[U_{\text{F}}^{(m,\langle e\rangle)}]^{t}=e^{-2it\theta_{em}}W_{i}^{(e)}$ ($t\in\mathbb{Z}$) hence $C_{ij}(t)=\bra{n}W_{i}^{(e)}(t)W_{j}^{(e)}\ket{n}=(-1)^{t}C_{ij}$. This period doubled long-ranged order is the spatio-temporal order defining TCs von Keyserlingk et al. (2016); Khemani et al. (2017, ). It also comes with spectral multiplets: if $U_{\text{F}}^{(m,\langle e\rangle)}\ket{n}=e^{-i\varepsilon_{n}}\ket{n}$, then, at the FP, $U_{\text{F}}^{(m,\langle e\rangle)}W_{i}^{(e)}\ket{n}=e^{2i\theta_{em}}e^{-i\varepsilon_{n}}W_{i}^{(e)}\ket{n}$; by $2\theta_{em}=\pi$, this is the familiar $\pi$ spectral pairing von Keyserlingk et al. (2016); Khemani et al. (2017, ).